OpenStudy (anonymous):

What percent of 60 is 12? Show your work

OpenStudy (anonymous):

x60=12 \[\LARGE \frac{x}{100}\cdot 60=12\] solve for x

OpenStudy (anonymous):

i took 60 * 12 % and got 7.2

OpenStudy (anonymous):

no ... solve the equation above I've written for x , and tell me what you get.

OpenStudy (anonymous):

\[\LARGE x\%60=12\] \[\LARGE \frac{x}{100}\cdot 60=12\] \[\LARGE \frac{x}{10}\cdot 6=12\] \[\LARGE 6x=120\] ... keep going when you have x=something that's it.

OpenStudy (anonymous):

@sexy1993 what did you get

OpenStudy (anonymous):

x=20

OpenStudy (anonymous):

that's correct. so What percent of 60 is 12? answer is 20%

OpenStudy (anonymous):